Nico is saving money for his college education. He invests some money at 8%, and $1200 less than that amount at 5%. The investments produced a total of $174 interest in 1 year. How much did he invest at each rate?
Answer by mananth(16949)

8%--------------- x
5% x-1200
Interest = 174
8%x+(x-1200)*5%= 174
multiply equation by 100
8x+5(x-1200)= 17400
8x+5x-6000=17400
13x= 17400+6000
13x= 23400
x=1800
Investment @ 8% = $1800
Investment at 5% = 1800-1200 = 600